OpenStudy (anonymous):

Ricardo ran the 400 meter 3 times. His fastest time was 54.3 sec. his slowest time was 56.1 sec. if is average time was 55.0 sec. what was his time for the third race?

OpenStudy (anonymous):

this is for my cousin he needs help

OpenStudy (ddcamp):

The average is the sum of the times divided by 3. If you multiply the average by 3, you get the total sum. \[3 \times average Time= Time1 +Time2+Time3 \\ 3 \times 55 = 54.3 + 56.1 + ?\]